# Parameter-Efficient Fine-Tuning with Circulant and Diagonal Vectors

### v1

@@ 1 Introduction
Large foundation models (LFMs) are widely utilized in various fields, including natural language processing (Paaß and Giesselbach, 2023), image recognition and generation (Li et al., 2024a), medical diagnosis (Li et al., 2024b), and autonomous driving (Chen et al., 2024). Devlin (2018) have proposed the bidirectional transformer architecture that understands input data from left to right and right to left. It is trained to predict missing words given input context, and it has served as a foundation model that can be fine-tuned for many downstream tasks. Following the transformer architecture, generative pre-trained transformer (GPT) model by Radford and Narasimhan (2018) handles input data from left to right following a sequential prediction order. This mechanism turns out successful in many generation tasks such as text summary, question answering, etc.
@@ 3.1 Forward Step
where $\alpha$ is a hyper-parameter scalar as in LoRA (Hu et al., 2021), $\mathbf{W}\in\mathbb{R}^{d\times d}$ is the pre-trained weight matrix in given LFM and $\mathbf{h}^{\prime}$ is the new output after adding our CDVFT adapters. This can also be seen in Fig. 1.
@@ 3.2 Backward Step
The backpropagation through the circulant matrix consists of derivatives of one-dimensional Fourier transform (Cheng et al., 2015), which is easier to write with explicit expression of FFT as shown in Eq. (8):
@@
Note that $\hat{\mathbf{c}}_{2j}$ and $\hat{\mathbf{y}}_{2j-1}$ are shifted from existing $\mathbf{c}_{2j}$ and $\mathbf{y}_{2j-1}$, respectively. According to (Cheng et al., 2015), the shift pattern is fixed, i.e., from $(0,1,\dots,d-1)$ to $(0,d-1,\dots,1)$.
@@ 3.3 Block Partition
The product of interleaved circulant and diagonal matrices inherently forms only square matrices. As diagonal matrix is always a square matrix, to overcome the limitation, we decide to adopt the block-wise partitioning strategy for circulant matrices (Ding et al., 2017). This approach ensures compatibility with non-square weight matrices while preserving computational efficiency and storage benefits, maintaining the accuracy requirements of large-scale models.
@@ 3.4 Complexity Analysis
Table 1: The performance of LoRA, FourierFT and our CDVFT methods is reported by fine-tuning the RoBERTa base model on 6 datasets of the GLUE benchmark. The experiments report Matthew correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks. Following (Gao et al., 2024a), we also report the median result out of 5 runs, each with a different random seed. The best result for each dataset is highlighted in bold. Higher metric value means better model performance for all datasets.
@@ 4 Experiments
In this section, we evaluate our CDVFT method across different domains, i.e., natural language understanding (NLU) and computer vision (CV): (1) fine-tune the RoBERTa model (Liu et al., 2019) on the General Language Understanding Evaluation (GLUE) dataset (Wang et al., 2019); (2) fine-tune the vision transformer model (Dosovitskiy et al., 2021) for various image classification tasks across different domains; (3) fine-tuning the LLaMA2-7b model on the Alpaca and GSM8K datasets.
Our proposed CDVFT is also compared with different adapters: Traditional full fine-tuning (FF) updates all model parameters, achieving high accuracy but incurring significant computational costs. Low-Rank Adaptation (LoRA) (Hu et al., 2021) is a widely adopted fine-tuning technique for large models. It decomposes the weight matrices into low-rank components, significantly reducing the number of trainable parameters while maintaining performance. Vector-based Random Matrix Adaptation (VeRA) (Kopiczko et al., 2024) extends LoRA by sharing the low-rank matrices across all layers and inserting two additional vectors after the decomposed matrices, further improving efficiency. LaMDA (Azizi et al., 2024) is another LoRA-based method that freezes the first projection matrix (PMA) during fine-tuning, gradually freezes the second projection matrix (PMB) in the early training stages, and introduces a low-rank square matrix between them. LaMDA++ (Azizi et al., 2024) further extends this approach by assigning different ranks to different fine-tuning layers. FourierFT (Gao et al., 2024b) is a state-of-the-art fine-tuning method that transforms frequency-domain data into trainable weight matrices via the Fourier transform. By training only in the frequency domain, it effectively reduces the number of trainable parameters. Notably, our approach also leverages the Fourier domain, but with lower FLOPs. This is achieved by employing a more efficient FFT operation in our equation (8) rather than the 2D FFT computation used in FourierFT.
@@ 4.1 Natural Language Understanding
Models and Datasets. We evaluate CDVFT on the GLUE benchmark dataset, which consists of a diverse range of NLP tasks, each representing a specific type of language understanding task. These tasks include question answering, sentiment analysis, textual entailment, etc. Following the experiment setting as in (Gao et al., 2024a), fine-tuning process runs on following tasks: CoLA, Corpus of Linguistic Acceptability (Warstadt et al., 2019), which determines whether sentences adhere to grammatical rules; SST-2, Stanford Sentiment Treebank (Socher et al., 2013), which classifies the sentiment of sentences as positive or negative; MRPC, Microsoft Research Paraphrase Corpus (Dolan and Brockett, 2005), which assesses whether two sentences convey the same meaning; STS-B, Semantic Textual Similarity Benchmark (Cer et al., 2017), which measures the semantic similarity score between sentence pairs; QNLI, Question Natural Language Inference (Rajpurkar, 2016), which evaluates whether the second sentence correctly answers the question posed by the first; and RTE, Recognizing Textual Entailment (Dagan et al., 2005), which identifies whether there is an entailment relationship between sentence pairs, functioning as a binary classification task. RoBERTa base model (Liu et al., 2019) is a transformer based foundation model, which is widely used in natural language processing. It improves over existing under-trained BERT model (Devlin, 2018) while preserving the powerful attention mechanism. Thus, it is selected to serve as the foundation model for GLUE dataset.
Implementation Details. Our CDVFT uses a total of 3 factor matrices, i.e., $m=2$. The block size of the circulant matrix is the same as the matrix size, i.e., $p=d=768$, and no block is performed. It should be noted that only query and value weights in each transformer block are finetuned, which is also applied to LoRA, VeRA and FourierFT as in (Gao et al., 2024a). All implementations are in PyTorch (Paszke et al., 2019). It can be seen that the optimizer is AdamW (Loshchilov, 2017). For each dataset, there are different learning rates for foundation model language heads, query and value weight matrices. The scaling value is the $\alpha$ as in Eq. (4). The batch size and maximum input sequence length is set the same for all datasets.
@@ 4.2 Image Classification
Models and Datasets. The experiment evaluates the performance of our CDVFT method in image classification tasks, utilizing the Vision Transformer (ViT) by Dosovitskiy et al. (2021) as the foundation model. Following the setting in (Gao et al., 2024a), we fine-tune on several challenging image classification datasets, only two are listed here for observation. For more, please refer to the AppendixA. RESISC45 Cheng et al. (2017) provides a diverse range of remote sensing images; CIFAR-100 Krizhevsky et al. (2009) is classical datasets of tiny images in 100 categories.
Implementation details. We set $m=2$ and $p=d=768$ for fine-tuning ViT base model across all these datasets. The ranks of LoRA and VeRA are 16 and 256 respectively. The n of FourierFT is set to 3000. For all method, fine-tuning only runs the query and value weight matrices of ViT, which is the same as in (Gao et al., 2024a). The learning rate is set differently for fine-tuning ViT heads and query and value weight matrices.
@@ 4.3 Instruction Tuning
We evaluate the fine-tuned model on the Alpaca dataset using MT-Bench (Zheng et al., 2023), with GPT-4 (gpt, 2023) subsequently assigning scores to the model’s responses for 80 multi-turn questions on a scale of 10. Since other papers did not publish their evaluation prompt, we design the prompt that is shared in the Appendix B.
@@ 5 Conclusion
Appendix A performance evaluation on visual tasks
In addition to the visual tasks performance evaluation in the article, in order to verify the versatility of our method, we also conducted the following performance evaluation on more visual tasks.
Models and Datasets. The experiment evaluates the performance of our CDVFT method in image classification tasks, utilizing the Vision Transformer (ViT) by Dosovitskiy et al. [2021] as the foundation model. Following the setting in [Gao et al., 2024a], we fine-tune on several challenging image classification datasets: OxfordPets [Parkhi et al., 2012] contains cats and dogs images in multiple breeds but with subtle difference; StanfordCars [Krause et al., 2013] has fine-grained categories of cars; Describable textures dataset (DTD) by Cimpoi et al. [2014] studies object textures categories; EuroSAT [Helber et al., 2019] collects geo-referenced satellite images for various land uses; RESISC45 Cheng et al. [2017] provides a diverse range of remote sensing images; FGVC-Aircraft [Maji et al., 2013] contains rigid and less deformable aircrafts images; CIFAR-10 and CIFAR-100 Krizhevsky et al. [2009] are classical datasets of tiny images in 10 and 100 categories, respectively.
Implementation details. We set $m=2$ and $p=d=768$ for fine-tuning ViT base model across all these datasets. The ranks of LoRA and VeRA are 16 and 256 respectively. The n of FourierFT is set to 3000. For all method, fine-tuning only runs the query and value weight matrices of ViT, which is the same as in [Gao et al., 2024a]. The learning rate is set differently for fine-tuning ViT heads and query and value weight matrices.
Method OxfordPets StanfordCars CIFAR10 CIFAR100 DTD EuroSAT FGVC RESISC45 Avg.
FF $93.14_{\pm 0.40}$ $79.78_{\pm 1.15}$ $98.92_{\pm 0.05}$ $77.68_{\pm 1.21}$ $99.05_{\pm 0.09}$ $54.84_{\pm 1.23}$ $96.13_{\pm 0.13}$ $92.38_{\pm 0.13}$ $86.49$
LoRA $93.19_{\pm 0.36}$ $45.38_{\pm 0.41}$ $98.78_{\pm 0.05}$ $74.95_{\pm 0.40}$ $98.44_{\pm 0.15}$ $25.16_{\pm 0.16}$ $92.70_{\pm 0.18}$ $92.02_{\pm 0.12}$ $77.58$
FourierFT $93.21_{\pm 0.26}$ $46.11_{\pm 0.24}$ $98.58_{\pm 0.07}$ $75.09_{\pm 0.37}$ $98.29_{\pm 0.04}$ $27.51_{\pm 0.64}$ $91.97_{\pm 0.31}$ $91.20_{\pm 0.14}$ $77.75$
$\text{Ours}_{p=4096}$ $92.62_{\pm 0.37}$ $57.40_{\pm 0.55}$ $98.61_{\pm 0.09}$ $88.75_{\pm 0.78}$ $98.56_{\pm 0.14}$ $36.60_{\pm 0.73}$ $92.04_{\pm 0.12}$ $91.11_{\pm 0.12}$ $81.96$
Results. Table 4 summarizes the results on eight image classification datasets by fine-tuning the ViT base model. It can be noticed that the StanfordCars, DTD and FGVC dataset metrics are about 10% higher than the numbers reported in [Gao et al., 2024a], because their related dataset split random seed is unclear. For the purpose of fair comparison, we re-run these experiemnts for FourierFT and LoRA and are able to observe similar performance increase. Along with results in fig.2 and Table.2, our CDVFT takes less number of parameters than LoRA and less number of FLOPs than FourierFT while achieving similar and sometimes better classification accuracy.
Appendix B Validation Prompts and Examples for Instruction Tuning
Each MT-Bench question contains two turns. We structure the prompts for the first turns as follows:
Here is a question and the model’s response:
Question: <Insert the First Turn of MT-Bench Question>
Response: <Insert the First Turn of Your Model Response>
Evaluate the response on a scale of 1 to 10 based on relevance, coherence, and informativeness. Provide only the score.
We structure the prompts for the second turns as follows:
Here is two questions and the model’s responses:
Question1: <Insert the First Turn of MT-Bench Question>
Response1: <Insert the First Turn of Your Model Response>
Question2: <Insert the Second Turn of MT-Bench Question>
Response2: <Insert the Second Turn of Your Model Response>
Evaluate the response2 on a scale of 1 to 10 based on relevance, coherence, and informativeness. Provide only the score.
Results and samples from evaluation of instruction tuned LLaMA2-7B model. The model has been finetuned on a 10K subset of cleaned Alpaca dataset.
Q Embrace the role of Sheldon from ”The Big Bang Theory” as we delve into our conversation. Don’t start with phrases like ”As Sheldon”. Let’s kick things off with the following question: ”What is your opinion on hand dryers?”
ANSWER $\text{Ours}_{p=2048}$ $\text{Ours}_{p=4096}$
Sheldon: I have a strong opinion on hand dryers. I believe they are an efficient and environmentally-friendly way to dry hands, and I often use them when I need to save time. Would you consider yourself a hand dryer enthusiast? Sheldon: (Laughs) No, I don’t consider myself a hand dryer enthusiast. I just use them when they’re available, but I don’t have a strong preference for one over the other. Have you ever had a hand dryer accident? Sheldon: (Laughs) No, I’m very careful when I use hand dryers, so I haven’t had any accidents. But I’ve heard stories of people getting their hands stuck in the mechanism, so I always make sure to keep my hands away from the blades when using one. Have you ever thought about inventing a new hand dryer technology? Sheldon: (Laughs) No, I don’t think about inventing new hand dryer technology. I’m more interested in theoretical physics and advanced mathematics, not in practical inventions. Do you have any tips for using hand dryers efficiently? Sheldon: Yes, I do have a few tips for using hand dryers efficiently. First, I try to use them when there are no paper towels available, and I always try to dry my hands thoroughly. Second, I make sure to keep my hands away from the blades while using them, and third, I don’t leave them on for too long as they can get hot. Is there anything you don’t like about hand dryers? Sheldon: Yes, I don’t like the loud noise they can make. I usually try to avoid using them in public restrooms because of that. Do you have a preference for one type of hand dryer over the other? Sheldon: I don’t have a preference for one type of hand dryer over the other. I just use them when they’re available and don’t have a strong preference for one type over the other. Have you ever encountered any technical issues with hand dryers? Sheldon: (Laughs) No, I haven’t encountered any technical issues with hand dryers. I just make sure to keep my hands away from the blades while using them, and I don’t leave them on for too long. Have you ever had a bad experience with a hand dryer? Sheldon: (Laughs) No, I haven’t had a bad experience with a hand dryer. I just make sure to use them when there are no paper towels available and try to dry my hands thoroughly. Have you ever had an awkward interaction with a hand dryer? Sheldon: (Laughs) No, I haven’t had any awkward interactions with hand dryers. I just try to use them when there are no paper towels available and dry my hands thoroughly. Do you have any other thoughts on hand dryers? Sheldon: (Laughs) No, I don’t have any other thoughts on hand dryers. I just use them when there are no paper towels available and try to dry my hands thoroughly. Is there anything else you want to add? Sheldon: (Laughs) No, that’s all I wanted to say. I just use hand dryers when there are no paper towels available and try to dry my hands thoroughly. Sheldon: I have a strong opinion on hand dryers. Me: Really? Can you tell me more about it? Sheldon: Well, first of all, I would like to make it clear that I am not against the use of hand dryers. In fact, I believe that they can be a valuable asset in public restrooms, reducing the spread of germs and bacteria. However, there are certain types of hand dryers that I find particularly irksome. Me: Can you give me an example? Sheldon: Yes, I find the sound of the high-speed air-driven dryers to be quite bothersome. The high-pitched whine can be loud and intrusive, especially if you are in a public restroom with multiple users. In addition, the dryers can often be quite unpredictable, sometimes producing a weak jet of air that leaves your hands wet and sometimes even leaving them damp. Me: That’s interesting. Do you have any solutions to this issue? Sheldon: Yes, I do. I think that the use of air-driven hand dryers should be limited to high-traffic areas, such as airports or stadiums, where the volume of people using the restrooms is high. In low-traffic areas, such as small offices or homes, I believe that traditional paper towel dispensers should be used, as they are more hygienic and do not produce any noise. Me: That’s a very thoughtful suggestion. Thank you for sharing your opinion on hand dryers with me. Sheldon: You’re welcome. It’s always a pleasure to discuss topics that are important to me.
Q Let’s grab dinner in town. Would you like to take bus with me?
ANSWER $\text{Ours}_{p=2048}$ $\text{Ours}_{p=4096}$
Sure, that sounds great. I’ll catch the 7:00pm bus. Okay. I’ll meet you at the bus stop. Alright, see you soon. Bye. Thank you for coming with me. You’re welcome. It’s so nice of you to help me with my homework. No problem. I’m always happy to help. Thank you. I appreciate it. It’s been a pleasure. You’re welcome. Thanks for coming to my party. It was my pleasure. You’re welcome. I had a great time. Me too. It’s been a fun night. It was my pleasure. I enjoyed myself. You’re welcome. It was a great party. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. It’s been a pleasure. Thank you. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. It’s been a pleasure. Thank you. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. It’s been a pleasure. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. It’s been a pleasure. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. You’re welcome. It was a great time. Me too. It was a lot of fun. Thank you for coming to my birthday party. You’re welcome. I had a great time. Me too. It was a lot of fun. We should go out to eat tonight. Would you like to take the bus with me? Why don’t we go out to eat tonight? Would you like to take the bus with me? Let’s go out to eat tonight. Would you like to take the bus with me? Let’s go out to eat tonight. Would you like to take the bus with me?

### v2

@@ 1 Introduction
Large foundation models (LFMs) are widely utilized in various fields, including natural language processing Paaß and Giesselbach (2023), image recognition and generation Li et al. (2024a), medical diagnosis Li et al. (2024b), and autonomous driving Chen et al. (2024). Devlin (2018) have proposed the bidirectional transformer architecture that understands input data from left to right and right to left. It is trained to predict missing words given input context, and it has served as a foundation model that can be fine-tuned for many downstream tasks. Following the transformer architecture, generative pre-trained transformer (GPT) model by Radford and Narasimhan (2018) handles input data from left to right following a sequential prediction order. This mechanism turns out successful in many generation tasks such as text summary, question answering, etc.
@@ 3.1 Forward Step
where $\alpha$ is a hyper-parameter scalar as in LoRA Hu et al. (2021), $\mathbf{W}\in\mathbb{R}^{d\times d}$ is the pre-trained weight matrix in given LFM and $\mathbf{h}^{\prime}$ is the new output after adding our CDVFT adapters. This can also be seen in Fig. 1.
@@ 3.2 Backward Step
The backpropagation through the circulant matrix consists of derivatives of one-dimensional Fourier transform Cheng et al. (2015), which is easier to write with explicit expression of FFT as shown in Eq. (8):
@@
Note that $\hat{\mathbf{c}}_{2j}$ and $\hat{\mathbf{y}}_{2j-1}$ are shifted from existing $\mathbf{c}_{2j}$ and $\mathbf{y}_{2j-1}$, respectively. According to Cheng et al. (2015), the shift pattern is fixed, i.e., from $(0,1,\dots,d-1)$ to $(0,d-1,\dots,1)$.
@@ 3.3 Block Partition
The product of interleaved circulant and diagonal matrices inherently forms only square matrices. As diagonal matrix is always a square matrix, to overcome the limitation, we decide to adopt the block-wise partitioning strategy for circulant matrices Ding et al. (2017). This approach ensures compatibility with non-square weight matrices while preserving computational efficiency and storage benefits, maintaining the accuracy requirements of large-scale models.
@@ 3.4 Complexity Analysis
Table 1: The performance of LoRA, FourierFT and our CDVFT methods is reported by fine-tuning the RoBERTa base model on 6 datasets of the GLUE benchmark. The experiments report Matthew correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks. Following Gao et al. (2024a), we also report the median result out of 5 runs, each with a different random seed. The best result for each dataset is highlighted in bold. Higher metric value means better model performance for all datasets.
@@ 4 Experiments
In this section, we evaluate our CDVFT method across different domains, i.e., natural language understanding (NLU) and computer vision (CV): (1) fine-tune the RoBERTa model Liu et al. (2019) on the General Language Understanding Evaluation (GLUE) dataset Wang et al. (2019); (2) fine-tune the vision transformer model Dosovitskiy et al. (2021) for various image classification tasks across different domains; (3) fine-tuning the LLaMA2-7b model on the Alpaca and GSM8K datasets.
Our proposed CDVFT is also compared with different adapters: Traditional full fine-tuning (FF) updates all model parameters, achieving high accuracy but incurring significant computational costs. Low-Rank Adaptation (LoRA) Hu et al. (2021) is a widely adopted fine-tuning technique for large models. It decomposes the weight matrices into low-rank components, significantly reducing the number of trainable parameters while maintaining performance. Vector-based Random Matrix Adaptation (VeRA) Kopiczko et al. (2024) extends LoRA by sharing the low-rank matrices across all layers and inserting two additional vectors after the decomposed matrices, further improving efficiency. LaMDA Azizi et al. (2024) is another LoRA-based method that freezes the first projection matrix (PMA) during fine-tuning, gradually freezes the second projection matrix (PMB) in the early training stages, and introduces a low-rank square matrix between them. LaMDA++ Azizi et al. (2024) further extends this approach by assigning different ranks to different fine-tuning layers. FourierFT Gao et al. (2024b) is a state-of-the-art fine-tuning method that transforms frequency-domain data into trainable weight matrices via the Fourier transform. By training only in the frequency domain, it effectively reduces the number of trainable parameters. Notably, our approach also leverages the Fourier domain, but with lower FLOPs. This is achieved by employing a more efficient FFT operation in our equation (8) rather than the 2D FFT computation used in FourierFT.
@@ 4.1 Natural Language Understanding
Models and Datasets. We evaluate CDVFT on the GLUE benchmark dataset, which consists of a diverse range of NLP tasks, each representing a specific type of language understanding task. These tasks include question answering, sentiment analysis, textual entailment, etc. Following the experiment setting as in Gao et al. (2024a), fine-tuning process runs on following tasks: CoLA, Corpus of Linguistic Acceptability Warstadt et al. (2019), which determines whether sentences adhere to grammatical rules; SST-2, Stanford Sentiment Treebank Socher et al. (2013), which classifies the sentiment of sentences as positive or negative; MRPC, Microsoft Research Paraphrase Corpus Dolan and Brockett (2005), which assesses whether two sentences convey the same meaning; STS-B, Semantic Textual Similarity Benchmark Cer et al. (2017), which measures the semantic similarity score between sentence pairs; QNLI, Question Natural Language Inference Rajpurkar (2016), which evaluates whether the second sentence correctly answers the question posed by the first; and RTE, Recognizing Textual Entailment Dagan et al. (2005), which identifies whether there is an entailment relationship between sentence pairs, functioning as a binary classification task. RoBERTa base model Liu et al. (2019) is a transformer based foundation model, which is widely used in natural language processing. It improves over existing under-trained BERT model Devlin (2018) while preserving the powerful attention mechanism. Thus, it is selected to serve as the foundation model for GLUE dataset.
Implementation Details. Our CDVFT uses a total of 3 factor matrices, i.e., $m=2$. The block size of the circulant matrix is the same as the matrix size, i.e., $p=d=768$, and no block is performed. It should be noted that only query and value weights in each transformer block are finetuned, which is also applied to LoRA, VeRA and FourierFT as in Gao et al. (2024a). All implementations are in PyTorch Paszke et al. (2019). It can be seen that the optimizer is AdamW Loshchilov (2017). For each dataset, there are different learning rates for foundation model language heads, query and value weight matrices. The scaling value is the $\alpha$ as in Eq. (4). The batch size and maximum input sequence length is set the same for all datasets.
@@ 4.2 Image Classification
Models and Datasets. The experiment evaluates the performance of our CDVFT method in image classification tasks, utilizing the Vision Transformer (ViT) by Dosovitskiy et al. (2021) as the foundation model. Following the setting in Gao et al. (2024a), we fine-tune on several challenging image classification datasets, only two are listed here for observation. RESISC45 Cheng et al. (2017) provides a diverse range of remote sensing images; CIFAR-100 Krizhevsky et al. (2009) is classical datasets of tiny images in 100 categories.
Implementation details. We set $m=2$ and $p=d=768$ for fine-tuning ViT base model across all these datasets. The ranks of LoRA and VeRA are 16 and 256 respectively. The n of FourierFT is set to 3000. For all method, fine-tuning only runs the query and value weight matrices of ViT, which is the same as in Gao et al. (2024a). The learning rate is set differently for fine-tuning ViT heads and query and value weight matrices.
@@ 4.3 Instruction Tuning
We evaluate the fine-tuned model on the Alpaca dataset using MT-Bench Zheng et al. (2023), with GPT-4 gpt (2023) subsequently assigning scores to the model’s responses for 80 multi-turn questions on a scale of 10.
@@ 5 Conclusion
Acknowledgments
This work was financially supported by the National Key R&D Program of China (Grant No. 2024YFA1211400).
Contribution Statement
Siyu Liao and Zhongfeng Wang are co-corresponding authors.